# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v2

@@ Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this letter, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative (phase rigidity) information supercurrents, analogous to a macroscopic superconducting phase governed by the Adler equation. The validity of this mapping is confirmed analytically and empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity, corresponding to the saddle-node bifurcation of phase-locking systems. These results define the new minimal chiral flocking model as a motile, disordered Josephson array, bridging active turbulence and superconductivity.
@@ I Introduction
We consider a recent, minimalist active matter turbulence model that is characterized by three simple rules, namely that its agents are (1) self-propelled, (2) polar, or flocking, favoring alignment, and (3) chiral, with intrinsic rotation [1, 2]. The model is unique in that it collapses these three properties into a single degree of freedom, phase, $\phi_{i}$,
@@
Of some importance, Refs. [1] and [2] both explicitly considered a renormalized fluid element in the model, the former by calculating the guiding centre location of each agent’s oscillatory motion, and the latter by the implementation of a low-pass filter scaled by the interaction correlation timescale. The papers motivated this step in terms of the Vlasov-Fokker-Planck equation [1] and shallow water theory [2], and the latter rigorously demonstrated that Eqs. (2) and (3) thereby lead to inviscid Euler turbulence in the form of an Onsager dipole.
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents, without a renormalized fluid element transformation, explicitly as a spatially extended, disordered Josephson junction array. The depinning transition between ’running’ and ’trapped’ states then generates a macroscopic synchronization stiffness. The resulting alignment torque between agents allows for an information supercurrent, the non-dissipative transmission of phase gradients across the ensemble.
@@ II Theoretical Foundation
Josephson junctions are essentially ensembles of quantum oscillators with a non-zero dispersion $\Delta\omega$ in their natural frequencies $\omega_{i}$, unavoidable due to fabrication heterogeneity [5, 6]. Josephson junctions synchronize (phase-lock) via a collective (global) coupling $K$ if the coupling strength is able to overcome the ensemble’s fabrication heterogeneity (frequency dispersion) $\Delta\omega$. Crucially, the coupling $K$ is naturally described with the Kuramoto-model [7], readily providing a theoretical foundation to understand the thermodynamics of our minimal model, which features a long-range, locally mediated Kuramoto interaction (Eqs. 2, 3).
@@
$v_{\text{det}}=\begin{cases}\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}\sim(I_{\text{eff}}-I_{c})^{1/2}&\text{ for }I_{\text{eff}}\geq I_{c}\\ \\ 0&\text{ for }I_{\text{eff}}<I_{c}\end{cases}$ (16)
@@
Figure 1: A simulated ensemble of polar chiral active matter agents. Panels a–d) show end-state snapshots of the entire simulation space, with agent point-cloud (panel a), topological charge (panel b), flow-lines of circulation in order $R^{2}$ (panel c), and agent slip velocity $v_{\text{slip}}$ (panel d), with respective colorscales where applicable. In panel a), trapped agents are shown in black while running agents are shown in red; corresponding to low and high $v_{\text{slip}}$ respectively, confirmed by panel d). Panel e) shows defect count as a function of normalized simulation time (left $y$-axis) with a $t^{-0.75}$ powerlaw fit shown with a green, dashed line, and the average potential barrier height $\Delta(t)$ (Eq. 18) in red (right $y$-axis, inverted). Panel f) shows a cross-correlation analysis of the latter two timeseries, for normalized times >5>5. Estimated confidence intervals for the cross-correlation are shown with dashed, black lines. Panels g–j) show inset $v_{\text{slip}}$ and trapped/running agent pointclouds, in rapid succession-snapshots, highlighting ongoing shock merger, agent thermalization, and subsequent re-organization. See Video S1 in the Supplementary Materials for a video of this simulation run.
III Results
@@
Figure 1g–j), showing four prior $v_{slip}$-snapshots of a limited portion of the simulation space, offer an instructional explanation of the model’s core thermodynamic mechanism. An ordered ($R\approx 1$) region shatters into phase slips (active turbulence), generating entropy; this is the shock merger mechanism identified by Ref. [2]. The agents inside this now-disordered patch are then quickly reorganized into the inertial fluid by the kinetic Turing instability [1]. It is this process that acts as a thermodynamic state preparation cycle [15], mimicking coherent microwave emissions in disordered Josephson arrays [14], effectively ’pumping’ the system from the high-entropy running state back into the low-entropy trapped state, thereby sustaining the Onsager dipole.
Figure 1e, f) demonstrate the foregoing dynamic in rigorous terms. First, we define the gap excess $\Delta(t)$, a metric derived from the potential barrier height of the overdamped Adler equation [11]. That is, we sum the potential heights for the trapped population,
@@
Panel e) shows topological defect count going down with a characteristic scaling of $t^{-0.75}$ [20, 21], identified by Ref. [2] to be caused by shock dynamics. At the same time, and with a strong negative correlation coefficient of 0.88-0.88 (Figure 1f), the gap excess $\Delta(t)$, or locking strength of the trapped agents [11], increase. The rise in $\Delta(t)$ during shock mergers confirms the topological heat pump mechanism: the annihilation of resistive defects ($\mathcal{U}_{B}\approx 0$) instantaneously throws agents into deep superconducting traps ($\mathcal{U}_{B}\gg 0$), actively lowering the system’s effective phase temperature.
Simultaneously, we observe in Figure 1(a,d) that agents trapped within the Onsager dipole’s vortex cores are continuously phase-slipping. These cores are identified by Ref. [2] as massless ($R\to 0$) sonic black holes. In this limit, the critical current Eq. (12) vanishes, reducing the Adler equation to its trivial kinematic limit ($\dot{\phi}_{i}\approx\omega_{i}$). Consequently, the vortex cores are purely dissipative singularities where the capacity to sustain supercurrents of phase rigidity has collapsed. This decoupling renders the cores unable to transmit phase stiffness or radiate coherent modes in the shallow water formalism, effectively silencing them within the inertial fluid.

### v3

@@ Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this article, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative phase rigidity supercurrents, a mapping we confirm empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity. Generalizing this framework to three dimensions ($S^{1}\to S^{2}$), we show that the polar alignment interaction is geometrically equivalent to the Gilbert damping term in spintronics. This constraint naturally generates inertial spin waves (ferromagnetic magnons) from the overdamped active bath, recovering the macroscopic transport predicted by Toner-Tu theory without invoking microscopic inertia. These results classify chiral active matter as a dissipative spintronic fluid, revealing that the transition to active turbulence is the emergence of a topological quasiparticle gas, where macroscopic Onsager dipoles acquire effective inertia from the Goldstone modes of the underlying broken symmetry.
@@ I Introduction
Active turbulence describes the chaotic spatiotemporal dynamics observed across a diverse spectrum of living and synthetic matter. This phenomenon is ubiquitous in non-equilibrium systems, having been rigorously characterized in the dense, swirling collective motion of bacterial swarms [1, 2], the restructuring of subcellular cytoskeletal extracts [3], and the engineered flows of synthetic colloidal flocks [4, 5]. The hydrodynamic conditions under which these behaviors emerge represents a fundamental challenge to statistical physics. Unlike classical inertial turbulence, which strictly requires high Reynolds numbers ($\mathcal{R}e$) to overcome viscosity and sustain energy cascades, active turbulence is overdamped ($\mathcal{R}e\approx 0$). In this viscous regime, where flows are expected to be laminar, chaotic transport is instead sustained by a distinct mechanism: the continuous, non-conservative injection of energy at the microscale, which drives the system far from equilibrium [6, 7].
A recent minimalist, two-dimensional chiral active matter model offers a solution to the paradox by empirically demonstrating that an overdamped chiral flock can drive a conservative inertial cascade that culminates in inviscid Euler turbulence, in the form of a stable, sloshing Onsager dipole in the dynamic agent ensemble, where topological defects interact long-range through the Kirchoff-Onsager hamiltonian [8] (Figure 1). The present article provides detailed theoretical backdrop for the findings in Ref. [8].
As we shall demonstrate, the agent inertia that is engineered into several successful hydrodynamic models of chiral flocks [9, 10, 11] can be replaced by a natural synchronization stiffness that mimicks an inertial flow, clearly observable in our minimalist model.
Figure 1: A simulated ensemble of polar chiral active matter agents. Panels a–d) show end-state snapshots of the entire simulation space, with agent point-cloud (panel a), topological charge (panel b), flow-lines of circulation in order $R^{2}$ (panel c), and agent slip velocity $v_{\text{slip}}$ (panel d), with respective colorscales where applicable. In panel a), trapped agents are shown in black while running agents are shown in red.
The model is characterized by three simple rules, namely that its agents are (1) self-propelled, (2) polar, or flocking, favoring alignment, and (3) chiral, with intrinsic rotation [12, 8]. The model is unique in that it collapses these three properties into a single degree of freedom, phase, $\phi_{i}$,
@@
Of some importance, Refs. [12] and [8] both explicitly considered a renormalized fluid element in the model, the former by calculating the guiding centre location of each agent’s oscillatory motion, and the latter by the implementation of a low-pass filter scaled by the interaction correlation timescale. The papers motivated this step in terms of the Vlasov-Fokker-Planck equation [12] and shallow water theory [8], and, as we have stated, the latter rigorously demonstrated that Eqs. (2) and (3) thereby lead to inviscid Euler turbulence in the form of an Onsager dipole.
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents as a spatially extended, disordered Josephson junction array. The depinning transition between ’running’ and ’trapped’ states then generates a macroscopic synchronization stiffness, or supercurrents of phase gradient information. By extending the model’s equations to three dimensions (3D), we eventually demonstrate that the rigidity supports Goldstone mode spin waves that map the two-dimensional (2D) model to the azimuthal precession of 3D active chiral agents. We shall contrast these ferromagnetic magnons against the second sound waves of the Toner-Tu field theory, delineating the universality of the Landau-Lifshitz-Gilbert equation.
@@ II Theoretical Foundation
In chiral flocks, motile agents that possess polar symmetry and seek to align, phenotypic heterogeneity [15] produces an intrinsic frequency dispersion among the flock constituents, in much the same way that semiconductor fabrication introduces a dispersion in the superconductor chips called Josephson junctions [16, 17]. Josephson junctions are essentially ensembles of quantum oscillators with a non-zero dispersion $\Delta\omega$ in their natural frequencies $\omega_{i}$, that synchronize (phase-lock) via a collective (global) coupling $K$ if the coupling strength is able to overcome the ensemble’s fabrication heterogeneity, or dispersion, $\Delta\omega$. Crucially, the coupling $K$ is naturally described with the Kuramoto-model [18], readily providing a theoretical foundation to understand the thermodynamics of our minimal model, which features a long-range, locally mediated Kuramoto interaction (Eqs. 2, 3).
@@
Figure 2: Panel a) shows defect count as a function of normalized simulation time (left $y$-axis) with a $t^{-0.75}$ powerlaw fit shown with a green, dashed line, and the average potential barrier height $\Delta(t)$ (Eq. 18) in red (right $y$-axis, inverted). Panel b) shows a cross-correlation analysis of the latter two timeseries, for normalized times >5>5. Estimated confidence intervals for the cross-correlation are shown with dashed, black lines. Panels c–j) show inset $v_{\text{slip}}$ and trapped/running agent pointclouds, data from the black square in Figure 1d, in rapid succession-snapshots, highlighting ongoing shock merger, agent thermalization, and subsequent re-organization. See Video S1 in the Supplementary Materials for a video of this simulation run.
@@
$v_{\text{det}}=\begin{cases}\sqrt{I_{\text{eff}}^{2}-I_{c}^{2}}\sim\sqrt{I_{\text{eff}}-I_{c}}&\text{ for }I_{\text{eff}}\geq I_{c}\\ \\ 0&\text{ for }I_{\text{eff}}<I_{c}\end{cases}$ (16)
@@
III Empirical Results
@@
Figure 2c–j), showing four prior $v_{slip}$-snapshots of a limited portion of the simulation space, offer an instructional explanation of the model’s core thermodynamic mechanism. An ordered ($R\approx 1$) region shatters into phase slips (active turbulence), generating entropy; this is the shock merger mechanism identified by Ref. [8]. The agents inside this now-disordered patch are then quickly reorganized into the inertial fluid by the kinetic Turing instability [12].
Figure 2a, b) demonstrate the foregoing dynamic in rigorous terms. First, we define the gap excess $\Delta(t)$, a metric derived from the potential barrier height of the overdamped Adler equation [22]. That is, we sum the potential heights for the trapped population,
@@
Figure 2a) shows topological defect count going down with a characteristic scaling of $t^{-0.75}$ [31, 32], identified by Ref. [8] to be caused by shock dynamics. At the same time, and with a strong negative correlation coefficient of 0.88-0.88 (Figure 2b), the gap excess $\Delta(t)$, or locking strength of the trapped agents [22], increase. The rise in $\Delta(t)$ during shock mergers confirms the topological heat pump mechanism: the annihilation of resistive defects ($\mathcal{U}_{B}\approx 0$) instantaneously throws agents into deep superconducting traps ($\mathcal{U}_{B}\gg 0$), actively lowering the system’s effective phase temperature.